A Level Mathematics 9709 Paper 51, Oct/Nov 2023

Question paper and mark scheme50 marks6 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Papers open in

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/51 · Oct/Nov 2023

Download PDF
1

The times taken by 120 children to complete a particular puzzle are represented in the cumulative frequency graph.

Figure 1 of 1 for question 1
a

Use the graph to estimate the interquartile range of the data.

[2]
b

35% of the children took longer than T seconds to complete the puzzle. Use the graph to estimate the value of T.

[2]
2

Hazeem repeatedly throws two ordinary fair 6-sided dice at the same time. On each occasion, the score is the sum of the two numbers that she obtains.

a

Find the probability that it takes exactly 5 throws of the two dice for Hazeem to obtain a score of 8 or more.

[2]
b

Find the probability that it takes no more than 4 throws of the two dice for Hazeem to obtain a score of 8 or more.

[2]
c

For 8 randomly chosen throws of the two dice, find the probability that Hazeem obtains a score of 8 or more on fewer than 3 occasions.

[3]
3

A farmer sells eggs. The weights, in grams, of the eggs can be modelled by a normal distribution with mean 80.5 and standard deviation 6.6. Eggs are classified as small, medium or large according to their weight. A small egg weighs less than 76 grams and 40% of the eggs are classified as medium.

a

Find the percentage of eggs that are classified as small.

[3]
b

Find the least possible weight of an egg classified as large.

[3]
c

150 of the eggs for sale last week were weighed. Use an approximation to find the probability that more than 68 of these eggs were classified as medium.

[5]
4

The times, to the nearest minute, of 150 athletes taking part in a charity run are recorded. The results are summarised in the table.

Time in minutes101−120121−130131−135136−145146−160
Frequency1848343218
a

Draw a histogram to represent this information.

[4]
b

Calculate estimates for the mean and standard deviation of the times taken by the athletes.

[5]
5

A red spinner has four sides labelled 1, 2, 3, 4. When the spinner is spun, the score is the number on the side on which it lands. The random variable X denotes this score. The probability distribution table for X is given below.

x1234
P(X=x)0.28p2p3p
a

Show that p=0.12.

[1]
b

A fair blue spinner and a fair green spinner each have four sides labelled 1, 2, 3, 4. All three spinners (red, blue and green) are spun at the same time. Find the probability that the sum of the three scores is 4 or less.

[3]
c

Find the probability that the product of the three scores is 4 or less given that X is odd.

[5]
6

In a restaurant, the tables are rectangular. Each table seats four people: two along each of the longer sides of the table (see diagram). Eight friends have booked two tables, X and Y. Rajid, Sue and Tan are three of these friends.

Figure 1 of 2 for question 6
Figure 2 of 2 for question 6
a

The eight friends will be divided into two groups of 4, one group for table X and one group for table Y. Find the number of ways in which this can be done if Rajid and Sue must sit at the same table as each other and Tan must sit at the other table.

[3]
b

When the friends arrive at the restaurant, Rajid and Sue now decide to sit at table X on the same side as each other. Tan decides that he does not mind at which table he sits. Find the number of different seating arrangements for the 8 friends.

[3]
c

As they leave the restaurant, the 8 friends stand in a line for a photograph. Find the number of different arrangements if Rajid and Sue stand next to each other, but neither is at an end of the line.

[4]

Mark scheme

9709/51 · Oct/Nov 2023

Download PDF

1a · 2 marks

  • M1

    [IQR =] 31 – 23.7

  • A1

    7.3

1b · 2 marks

  • B1

    [65% of 120 = ]78

  • B1

    28.5

2a · 2 marks

  • M1

    (2136)4(1536)

  • A1

    =12005248832,0.0482

2b · 2 marks

  • M1

    [P(X⩽4)=]1−(2136)4 or [P(X⩽4)=]1536+1536×2136+1536×(2136)2+1536×(2136)3

  • A1

    =1833520736,0.884

2c · 3 marks

  • M1

    One term 8Cx(q)x(1−q)8−x,0<q<1,x≠0,8.

  • A1

    Correct expression, accept unsimplified, no terms omitted leading to final answer. FT only with unsimplified expression.

  • B1

    = 0.282

3a · 3 marks

  • M1

    [P(X<76)=]P(Z<76−80.56.6)

  • M1

    [=Φ(−0.6818)=1−Φ(0.6818)=] 1−0.7524=0.2476

  • A1

    24.8%

3b · 3 marks

  • B1

    0.378⩽z<0.3791 or −0.3791<z⩽−0.378 seen.

  • M1

    x−80.56.6=0.378

  • A1

    x=83[.0]

3c · 5 marks

  • B1

    Mean=150×0.4=60 Var=150×0.4×0.6=36

  • M1

    P(X>68)=P(Z>68.5−6036)

  • M1

    Using continuity correction 67.5 or 68.5 in their standardisation formula.

  • M1

    P(Z>1.417)=1−Φ(1.417) [=1−0.9217]

  • A1

    0.0783

4a · 4 marks

  • M1

    At least 4 frequency densities calculated by fcw, e.g. 1820 (condone fcw±0.5 if unsimplified). Accept unsimplified, may be read from graph using their scale, no lower than 1cm = 1 fd.

  • A1

    All bar heights correct on graph (no FT), using their suitable linear scale with at least 3 values indicated, no lower than 1cm = 1 fd.

  • B1

    Bar ends at 120.5, 130.5, 135.5, 145.5, 160.5. 5 bars drawn with a horizontal linear scale, no lower than 1 cm = 10min, with at least 3 values indicated. 100 ⩽ horizontal scale ⩽ 160.

  • B1

    Axes labelled frequency density (fd), time (t) and minutes (min, m) oe, or an appropriate title. (Axes may be reversed).

4b · 5 marks

  • B1

    [Midpoints ] 110.5 125.5 133 140.5 153

  • M1

    Mean=18×110.5+48×125.5+34×133+32×140.5+18×153150=1989+6024+4522+4496+2754150

  • A1

    = 131.9

  • M1

    Variance=18×110.52+48×125.52+34×1332+32×140.52+18×1532150−(their 131.9)2

  • A1

    [Standard deviation =] 11.7

5a · 1 mark

  • B1

    0.28+6p=1, p=0.12

5b · 3 marks

  • B1

    Correct probability for 1 identified scenario, accept unsimplified, www.

  • M1

    Add values of 4 correct scenarios, may be implied by correct unsimplified expressions. No incorrect/repeated scenarios.

  • A1

    0.06

5c · 5 marks

  • B1

    [P(X is odd) = 0.28 + 2×0.12 or 0.24 ]= 0.52[0]

  • M1

    Values of at least 5 identified correct scenarios added, accept unsimplified, condone incorrect scenarios in calculation.

  • M1

    [P(product of 3 scores ⩽ 4 ∩ X is odd) = ] 0.28×(0.25)2×8+0.24×(0.25)2

  • M1

    [P(product of 3 scores ⩽4∣X is odd)=P(product of 3 scores ⩽4∩X is odd)P(X is odd)=]0.1550.52

  • A1

    =0.298,155520,31104

6a · 3 marks

  • M1

    5C2×r,r=positive integer, 1 implied, no addition.

  • M1

    s×2,s=5C2 or 5P2 or if 5C2 or 5P2 not present, s=a single integer>1 or t!×2,2⩽t⩽8, no other terms.

  • A1

    20

6b · 3 marks

  • M1

    Method 1: 6C2×2×2×2×t,t=positive integer⩾1 6P2×2×2×t,t=positive integer⩾1 Method 2: 6!×v,v=positive integer⩾1

  • M1

    Method 1: u×4!,u=positive integer>1 Method 2: w×2×2,w=positive integer>1 condone w×4,w=positive integer>1

  • A1

    2880

6c · 4 marks

  • M1

    Method 1: 7! × 2 – a, a = positive integer > 1. Method 2: 6! × e × f, e, f = positive integers ⩾ 1. Method 3: ⁶P₂ × e × f, e, f = positive integers ⩾ 1. Method 4: ⁶Pₙ × a × (6−n)!, a = positive integer, 1 ⩽ n ⩽ 5 seen once.

  • M1

    Method 1: b – 6! × 4, b = positive integer > 2880. Method 2: 6! × 2 × f, f = positive integer ⩾ 1. If 5! Used, SC B1 5! × 2 × f, f = positive integer > 1. Method 3: ⁶P₂ × 5! × f, f = positive integer ⩾ 1. Condone ⁶C₂ × 5! × f, f = positive integer ⩾ 1. Method 4: ⁶Pₙ × 2 × (6−n)!, a = positive integer, 1 ⩽ n ⩽ 5 seen at least 3 times in identified scenarios.

  • M1

    Method 1: 7! × c – 6! × d, c = 1,2 and d = 1, 4. Method 2: 6! × e × 5, e = positive integer ⩾ 1. Method 3: ⁶P₂ × e × 2, e = positive integer ⩾ 1. Condone ⁶C₂ × e × 2, e = positive integer ⩾ 1. Method 4: Add 5 values of appropriate scenarios only. No additional, incorrect or repeated scenarios. Accept unsimplified.

  • A1

    7200

Questions

  1. 1
    Question 14 marks
    Mark scheme ›
  2. 2
    Question 27 marks
    Mark scheme ›
  3. 3
    Question 311 marks
    Mark scheme ›
  4. 4
    Question 49 marks
    Mark scheme ›
  5. 5
    Question 59 marks
    Mark scheme ›
  6. 6
    Question 610 marks
    Mark scheme ›

More 9709 papers

Other papers from Oct/Nov 2023

Solve it on ExamPalStart